OpenStudy (anonymous):

Help!!! I also have 2 more!! please Elsa knows that for every bushel of apples she sells she earns $9 and for every basket of pears she earns $11. She uses the expression 9a + 11p to keep track of her earnings. Part A: Identify the coefficients and variables in the expression. (3 points) Part B: How many terms are in the expression, what are they, and how do you know? (4 points) Part C: Which term in the expression shows the total earned from selling baskets of pears?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

Jolie can read around 14 sentences per minute of her current novel. She wants to finish her reading before the movie comes out next week. If she has 17 pages left to read and each page has about 120 sentences, how many hours should Jolie set aside to complete her reading? Show your work to receive full credit. (10 points)

OpenStudy (anonymous):

Donna's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $50 a day for food and lodging and $0.60 for each mile traveled. Donna drove 200 miles and was reimbursed $1,620. Part A: Create an equation that will determine the number of days x on the trip. (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did Donna spend on this trip? (1 point

OpenStudy (word2):

okay 1 at a time , for the first one . the coefficients are 9 and 11, and the variables are a and p there are 2 terms 9(a) and 11(p) you can tell because they have a + in between and you can use 11p to find the total from selling pears

OpenStudy (anonymous):

@Elsa213

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@Hero @haleyelizabeth2017

OpenStudy (anonymous):

Number of sentences left to read is 17x120=2040 Number of minutes it will take=2040/14 =145.71minutes 145.71 minutes= 2 hours 25.71minutes Jolie needs to set aside 2 hours 26minutes is this right?

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Coefficients are 9 and 11, and the variables are a and p there are 2 terms 9a and 11p you can tell because they have a + inbetween and you can use 11p to find the total from selling pears. Is this right???

OpenStudy (johnweldon1993):

Both look good! What about the third?

OpenStudy (anonymous):

A) $50.00x + 200($0.60) = $1,620 B) $50x + $120 = $1620 $50x = $1500 x = 30 C) 30 days

OpenStudy (anonymous):

@johnweldon

OpenStudy (anonymous):

@johnweldon1993